Robinson Hardware is adding a new product line that will require an investment of $1,500,000. Managers estimate that this investment will have a​ 10-year life and generate net cash inflows of $320,000 the first​ year, $265,000 the second​ year, and $230,000 each year thereafter for eight years. The investment has no residual value. Compute the ARR for the investment.

​First, enter the​ formula, then compute the ARR of the new product line.
Answer #1

Solution:

Annual depreciation = $1,500,000 / 10 = $150,000

Average annual cash inflows = ($320,000 + $265,000 + $230,000*8) / 10 = $242,500

Average annual income = $242,500 - $150,000 = $92,500

Average investment = (Cost + Residual value) / 2 = ($1,500,000 +0)/2 = $750,000

ARR = Average annual income / Average investment = $92,500 / $750,000 = 12.33%